# Path Eccentricity and Forbidden Induced Subgraphs

### v1

@@ Abstract
[…]
@@ 1 Introduction
Here, a claw is the complete bipartite graph $K_{1,3}$, the net is the graph obtained by appending a pendant edge to each vertex of the complete graph $K_{3}$, and, given a set $\mathcal{F}$ of graphs, a graph $G$ is said to be $\mathcal{F}$-free if $G$ does not have an induced subgraph isomorphic to any member of $\mathcal{F}$. If =$\mathcal{F}=\{F\}$ for some graph $F$, we also say that $G$ is $F$-free.
The result of Theorem 1.1 is tight, since the claw and the net do not have a Hamiltonian path. Let us remark that, in general, if $G$ is a graph that has a Hamiltonian path and $H$ is a connected induced subgraph of $G$, then $H$ may fail to have a Hamiltonian path. For example, if $H$ is any $n$-vertex graph and $G$ is a graph obtained from $H$ by adding to it $n$ new vertices fully adjacent to the vertices of $H$, then $G$ has a Hamiltonian path, even though $H$ may fail to have one. In particular, this construction could be applied to the the claw and the net.
If we instead require that all connected induced subgraph have a Hamiltonian path, Theorem 1.1 implies the following characterization.
@@
In 2023, Gómez and Gutiérrez [GG23] generalized the concept of Hamiltonian paths by defining the notion of path eccentricity of graphs. Given a connected graph $G$ and a path $P$ in $G$, the eccentricity of $P$ is the maximum distance from a vertex in $G$ to a vertex in $P$. Hence, in particular, a path $P$ has eccentricity 00 if and only if $P$ is a Hamiltonian path, and $P$ has eccentricity at most $1$ if and only if $P$ is dominating, that is, every vertex not in the path has a neighbor on the path. More generally, given an integer $k\geqslant 0$, a path $P$ is said to be $k$-dominating if it has eccentricity at most $k$. The path eccentricity of a connected graph $G$ is denoted by $\operatorname{\mathbf{pe}}(G)$ and defined as the minimum eccentricity of a path in $G$. Hence, $\operatorname{\mathbf{pe}}(G)=0$ if and only if $G$ has a Hamiltonian path, and $\operatorname{\mathbf{pe}}(G)\leqslant 1$ if and only if $G$ has a dominating path. It is not difficult to see that a graph $G$ has a dominating path if and only if $G$ has a spanning caterpillar, that is, a spanning subgraph that is a caterpillar (a tree possessing a dominating path). Hence, understanding graphs that admit dominating paths is motivated by their applications to the graph burning and a variant of the cops and robber game. More precisely:
@@
  • If a graph $G$ admits a spanning caterpillar $T$, then the minimum number of cops sufficient for the cop player to win on the graph $G$ against the robber moving at the speed of $s$ is at most $ps$ if $G$ is a subgraph of the graph obtained from $T$ by adding an edge between each pair of vertices that are at distance at most $p$ in $T$ (see Fomin et al. [FGK+10, Lemma 4]).
Since adding a universal vertex to any connected graph $H$ results in a graph $G$ with path eccentricity at most $1$, path eccentricity is not monotone under vertex deletion. Nevertheless, and in line with the above discussion preceding Corollary 1.2, which characterizes graphs each connected induced subgraph of which has path eccentricity 00, analogous questions can be addressed for higher values of path eccentricity.
@@
For each fixed $k\geqslant 1$, there are two obstructions to path eccentricity less than $k$ that can be obtained from the claw and the net, as follows. Given a positive integer $k$, let us denote by $S_{k}$ the $k$-subdivided claw, that is, the graph obtained from the claw by replacing each of its edges with a path of length $k$ (in particular, $S_{1}$ is the claw). Furthermore, we denote by $T_{k}$ the $k$-subdivided net, that is, the graph obtained by appending a path of length $k$ to each vertex of the complete graph $K_{3}$ (in particular, $T_{1}$ is the net). See Figure 1 for an example.
@@
It is easy to observe that $\operatorname{\mathbf{pe}}(S_{k})=\operatorname{\mathbf{pe}}(T_{k})=k$ for each positive integer $k$. Consequently, if $G$ is a graph such that each connected induced subgraph of $G$ has path eccentricity less than $k$, then $G$ is $\{S_{k},T_{k}\}$-free. As the main result of this paper, we show that this necessary condition is in fact also sufficient. We thus obtain the following unifying theorem, valid for all positive integers $k$ (and thus capturing also Corollary 1.2, which corresponds to the case $k=1$).
@@
The case $k=2$ of Theorem 1.4 characterizes graphs every connected induced subgraph of which has a dominating path, and implies the following.
@@
Corollary 1.5 unifies and generalizes several results from the literature regarding sufficient conditions for the existence of a dominating path:
@@
    A graph $G$ is said to be AT-free if it does not have an independent set $I$ such that $|I|=3$ and for every vertex $v\in I$, the remaining two vertices in $I$ are in the same component of $G-N[v]$. Every connected AT-free graph has a dominating path (see Corneil, Olariu, and Stewart [COS97], as well as [COS95]).
@@
    A graph $G$ is biconvex if it is bipartite, with parts $A$ and $B$ that can each be linearly ordered so that for each vertex $v$ of $G$, the neighborhood of $v$ in the part not containing $v$ forms a consecutive segment of vertices with respect to the linear ordering. Gómez and Gutiérrez [GG23] and independently Antony et al. [ADG+24] proved that every connected biconvex graph has a dominating path (or, equivalently, a spanning caterpillar).
Indeed, each of the previously stated cases corresponds to a subclass of the class of $\{S_{2},T_{2}\}$-free graphs, hence, Corollary 1.5 implies all these results.
Further consequences of Theorem 1.4 are related to a recent work of Bastide, Hilaire, and Robinson [BHR25]. The authors explored a relation between the path eccentricity of a graph and a certain property of its adjacency matrix (or equivalently on the neighborhoods of its vertices). They introduced the partially augmented consecutive ones property (denoted *-C1P for short): a graph has the *-C1P if there exists an ordering of its vertices sending the open or closed neighborhood of every vertex to a consecutive set. The authors showed that graphs with the *-C1P have path eccentricity at most 2, and conjectured that they actually have path eccentricity at most 1. We show that Theorem 1.4 resolves the conjecture.
@@
Another result from [BHR25] is related to the following generalization of AT-freeness. A graph $G$ is said to be $k$-AT-free if $G$ does not have an independent set $I$ such that $|I|=3$ and for every vertex $v\in I$, the remaining two vertices in $I$ are in the same component of the graph $G-\{u\colon u$ is at distance at most $k$ of $v\}$. In particular, a graph is AT-free if and only if it is $1$-AT-free. In [BHR25], the authors proved that for every $k\geqslant 1$, if a graph $G$ is $k$-AT-free, then $\operatorname{\mathbf{pe}}(G)\leqslant k$. This result follows from Theorem 1.4. Indeed, it is easy to check that for every $k\geqslant 2$, the graphs $S_{k}$ and $T_{k-1}$ are not $(k-1)$-AT-free. Thus, if a graph $G$ is $(k-1)$-AT-free, then $G$ is $\{S_{k},T_{k}\}$-free, which implies by Theorem 1.4, that $\operatorname{\mathbf{pe}}(G)<k$.
For the case where a single induced subgraph is excluded, Theorem 1.4 leads to the following characterization. The symbol $+$ denotes the disjoint union of graphs, and for an integer $k\geqslant 0$ and a graph $H$, we denote by $kH$ the disjoint union of $k$ copies of $H$.
@@
In the case of $3P_{2}$-free graphs, we obtain a much stronger property. We can find a path with eccentricity at most $1$ that is a longest path in the graph. While the proof of this result implies that we can transform a longest path into a longest path that is dominating in polynomial time, it is NP-hard to find a longest path in $3P_{2}$-free graphs. In fact, determining whether a given graph has a Hamiltonian path is NP-complete in the class of split graphs [Mül96] that are even $2P_{2}$-free.
@@
Every connected $3P_{2}$-free graph $G$ has a longest path that is dominating. Moreover, we can find a (not necessarily longest) dominating path in $G$ in time $\mathcal{O}(n^{2}(n+m))$, where $n$ is the number of vertices and $m$ is the number of edges in $G$.
@@
This is a very fine line, as this property is completely violated in the case of $\{S_{2},T_{2}\}$-free graphs or even $\{P_{5},4P_{2}\}$-free graphs; see Figure 2. It also cannot be generalized to higher eccentricity; see for example a $\{3P_{3},P_{7}+P_{2}\}$-free graph on Figure 3. Moreover, in $3P_{2}$-free graph there might exist a longest path that is not dominating, as depicted in Figure 4.
Figure 3: An example of a $\{3P_{3},P_{7}+P_{2}\}$-free graph no longest path of which is $2$-dominating. The vertices depicted in orange need to belong to every $2$-dominating path; however, no longest path contains the top orange vertex.
@@
While Theorem 1.8 actually provides us with a polynomial-time algorithm that finds a dominating path, the algorithm does not necessarily find a longest path that is dominating, unless we are given another longest path on the input.
@@
In Section 2, we give an overview of known results on structural domination, including a key result of Bacsó and Tuza [BT12] that we use in the proof of our main theorem (Theorem 1.4). We prove Theorems 1.4 and 1.7 in Section 3. We prove Theorem 1.6 in Section 4. We prove Theorem 1.8 in Section 5. We conclude the paper with some final remarks and open questions in Section 6.
@@ 2 Background on Structural Domination
Theorem 1.4 is closely related to the framework of structural domination. Given a class $\mathcal{D}$ of connected graphs, let us say that a graph $G$ is hereditarily dominated by $\mathcal{D}$ if each connected induced subgraph of $G$ contains a dominating set that induces a graph from $\mathcal{D}$. Bacsó, Michalak, and Tuza (see [BMT05]) gave forbidden induced subgraph characterizations of graphs hereditarily dominated by classes of complete bipartite graphs, stars, connected bipartite graphs, and complete $k$-partite graphs. Similar results were obtained by Michalak (see [Mic07]), for graphs hereditarily dominated by classes of cycles and paths, paths, trees, and trees with bounded diameter. Camby and Schaudt (see [CS16]) proved that for every $k\geqslant 4$, a graph $G$ is $P_{k}$-free if and only if it is hereditarily dominated by the class consisting of all connected $P_{k-2}$-free graphs along with the $k$-vertex cycle.
The most general result along these lines was obtained independently by Bacsó and Tuza (see [Tuz08, Bac09]). To state the result, we need one more definition. Given a connected graph $H$, the leaf graph of $H$ is the graph $H^{\text{\textleaf}}$ obtained from $H$ by attaching a leaf to each vertex $v\in V(H)$ that is not a cut-vertex (that is, the graph $H-v$ is connected). By definition, these new leaves are pairwise distinct and non-adjacent.
@@
The main result of [Tuz08, Bac09] is even more general, without the restriction that $\mathcal{D}$ contains the class of all paths; but since we do not need this result, we omit the precise statement.
A generalization of Theorem 2.1 to domination at distance was proved by Bacsó and Tuza (see [BT12]). A distance-$k$ dominating set in a graph $G$ is a set $D\subseteq V(G)$ such that every vertex in $G$ is at distance at most $k$ from some vertex in $D$. If, in addition, the graph $G[D]$ is connected, then $D$ is said to be a connected distance-$k$ dominating set. A graph $G$ is said to be hereditarily $k$-dominated by a class $\mathcal{D}$ of connected graphs if each connected induced subgraph of $G$ admits a distance-$k$ dominating set that induces a graph from $\mathcal{D}$. Given a connected graph $H$, the $k$-leaf graph of $H$ is the graph $H_{k}^{\text{\textleaf}}$ obtained from $H$ by attaching a pendant path of length $k$ to each vertex $v\in V(H)$ that is not a cut-vertex. Bacsó and Tuza (see [BT12]) generalized Theorem 2.1 as follows.
@@
Note that Theorem 2.1 is a special case of Theorem 2.2 for the case $k=1$.
3 Proof of General Theorem (Theorem 1.4)
@@
Fix an integer $k\geqslant 1$. Let $G$ be a graph such that every connected induced subgraph $H$ of $G$ satisfies $\operatorname{\mathbf{pe}}(H)<k$. To show that $G$ is $\{S_{k},T_{k}\}$-free, it suffices to show that $\operatorname{\mathbf{pe}}(S_{k})\geqslant k$ and $\operatorname{\mathbf{pe}}(T_{k})\geqslant k$. Let $H$ be a graph isomorphic to either $S_{k}$ or $T_{k}$ and let $S$ be any path in $H$. We claim that $S$ has eccentricity at least $k$. Suppose not. Let $C$ be the set of vertices of degree $3$ in $H$. Then, the graph $H-V(C)$ consists of three disjoint paths $P$, $Q$, and $R$, each with $k$ vertices. Moreover, each of the paths $P$, $Q$, and $R$ contains a vertex at distance $k$ from $V(C)$. Since the eccentricity of $S$ is less than $k$, each of the paths $P$, $Q$, and $R$ contains a vertex of $S$. Since $S$ has at most two endpoints, at most two of the paths $P$, $Q$, and $R$ contain an endpoint of $P$; hence, we may assume without loss of generality that $P$ contains an internal vertex of $S$ but no endpoint of $S$. However, this implies that two distinct edges of $S$ connect $P$ with the rest of $H$, a contradiction. This shows that $\operatorname{\mathbf{pe}}(H)\geqslant k$, as claimed.
It remains to prove that if $G$ is $\{S_{k},T_{k}\}$-free, then every connected induced subgraph $H$ of $G$ has path eccentricity less than $k$, or, equivalently, that every connected $\{S_{k},T_{k}\}$-free graph has path eccentricity less than $k$. For $k=1$, the statement coincides with that of Theorem 1.1. Suppose now that $k\geqslant 2$. Let =$\mathcal{F}=\{S_{1},T_{1}\}$ and let $\mathcal{D}$ be the class of all connected $\mathcal{F}$-free graphs. Note that $\mathcal{F}_{k-1}^{\text{\textleaf}}=\{S_{k},T_{k}\}$. Consequently, since $\mathcal{D}$ satisfies the hypothesis of Theorem 2.2, it follows that every $\{S_{k},T_{k}\}$-free graph $G$ is hereditarily $(k-1)$-dominated by $\mathcal{D}$. Let $H$ be a connected induced subgraph of $G$. Then, since $G$ is hereditarily $(k-1)$-dominated by $\mathcal{D}$, there exists a distance-$(k-1)$ dominating set $D$ in $H$ such that the induced subgraph $H[D]$ belongs to $\mathcal{D}$. By Theorem 1.1, the graph $H[D]$ admits a Hamiltonian path $P$. It follows that $P$ is a path in $H$ with eccentricity less than $k$, implying that $\operatorname{\mathbf{pe}}(H)<k$. ∎
Our proof of Corollary 1.7 will be based on the following.

### v2

# Jana Masaříková

@@ Abstract
The path eccentricity of a connected graph $G$ is the minimum integer $k$ such that $G$ has a path such that every vertex is at distance at most $k$ from the path. A result of Duffus, Jacobson, and Gould from 1981 states that every connected $\{\text{claw},\text{net}\}$-free graph $G$ has a Hamiltonian path, that is, $G$ has path eccentricity 0. Several more recent works identified various classes of connected graphs with path eccentricity at most $1$, or, equivalently, graphs having a spanning caterpillar, including connected $P_{5}$-free graphs, AT-free graphs, and biconvex graphs. Generalizing all these results, we apply the work on structural distance domination of Bacsó and Tuza [Discrete Math., 2012] and characterize, for every positive integer $k$, graphs such that every connected induced subgraph has path eccentricity less than $k$. More specifically, we show that every connected $\{S_{k},T_{k}\}$-free graph has a path eccentricity less than $k$, where $S_{k}$ and $T_{k}$ are two specific graphs of path eccentricity $k$ (a subdivided claw and the line graph of such a graph). As a consequence, every connected $H$-free graph has path eccentricity less than $k$ if and only if $H$ is an induced subgraph of $3P_{k}$ or $P_{2k+1}+P_{k-1}$. For such cases, we also provide a robust polynomial-time algorithm that finds a path witnessing the upper bound on the path eccentricity. Our main result also answers an open question of Bastide, Hilaire, and Robinson [Discrete Math., 2025].
@@ 1 Introduction
Here, a claw is the complete bipartite graph $K_{1,3}$, the net is the graph obtained by appending a pendant edge to each vertex of the complete graph $K_{3}$, and, given a set $\mathcal{F}$ of graphs, a graph $G$ is said to be $\mathcal{F}$-free if $G$ does not have an induced subgraph isomorphic to any member of $\mathcal{F}$. If $\mathcal{F}=\{F\}$ for some graph $F$, we also say that $G$ is $F$-free.
The result of Theorem˜1.1 is tight, since the claw and the net do not have a Hamiltonian path. Let us remark that, in general, if $G$ is a graph that has a Hamiltonian path and $H$ is a connected induced subgraph of $G$, then $H$ may fail to have a Hamiltonian path. For example, if $H$ is any $n$-vertex graph and $G$ is a graph obtained from $H$ by adding to it $n$ new vertices fully adjacent to the vertices of $H$, then $G$ has a Hamiltonian path, even though $H$ may fail to have one. In particular, this construction could be applied to the claw and the net.
If we instead require that all connected induced subgraphs have a Hamiltonian path, Theorem˜1.1 implies the following characterization.
@@
In 2023, Gómez and Gutiérrez [GG23] generalized the concept of Hamiltonian paths by defining the notion of path eccentricity of graphs. Given a connected graph $G$ and a path $P$ in $G$, the eccentricity of $P$ is the maximum distance from a vertex in $G$ to a vertex in $P$. Hence, in particular, a path $P$ has eccentricity 0 if and only if $P$ is a Hamiltonian path, and $P$ has eccentricity at most $1$ if and only if $P$ is dominating, that is, every vertex not in the path has a neighbor on the path. More generally, given an integer $k\geqslant 0$, a path $P$ is said to be $k$-dominating if it has eccentricity at most $k$. The path eccentricity of a connected graph $G$ is denoted by $\operatorname{\mathbf{pe}}(G)$ and defined as the minimum eccentricity of a path in $G$. In particular, $\operatorname{\mathbf{pe}}(G)=0$ if and only if $G$ has a Hamiltonian path, and $\operatorname{\mathbf{pe}}(G)\leqslant 1$ if and only if $G$ has a dominating path. It is not difficult to see that a graph $G$ has a dominating path if and only if $G$ has a spanning caterpillar, that is, a spanning subgraph that is a caterpillar (a tree possessing a dominating path). Hence, understanding graphs that admit dominating paths is motivated by their applications to graph burning and a variant of the cops and robber game. More precisely:
@@
  • If a graph $G$ admits a spanning caterpillar $T$, then the minimum number of cops sufficient for the cop player to win on the graph $G$ against the robber moving at the speed of $s$ is at most $ps$ if $G$ is a subgraph of the graph obtained from $T$ by adding an edge between each pair of vertices that are at distance at most $p$ in $T$ (see Fomin, Golovach, Kratochvíl, Nisse, and Suchan [FGK+10, Lemma 4]).
Since adding a universal vertex to any connected graph $H$ results in a graph $G$ with path eccentricity at most $1$, path eccentricity is not monotone under vertex deletion. Nevertheless, and in line with the above discussion preceding Corollary˜1.2, which characterizes graphs each connected induced subgraph of which has path eccentricity 0, analogous questions can be addressed for higher values of path eccentricity.
@@
For each fixed $k\geqslant 1$, there are two obstructions to path eccentricity less than $k$ that can be obtained from the claw and the net, as follows. Given a positive integer $k$, let us denote by $S_{k}$ the $k$-subdivided claw, that is, the graph obtained from the claw by replacing each of its edges with a path of length $k$, where the length of a path is the number of its edges (in particular, $S_{1}$ is the claw). Furthermore, we denote by $T_{k}$ the $k$-subdivided net, that is, the graph obtained by appending a path of length $k$ to each vertex of the complete graph $K_{3}$ (in particular, $T_{1}$ is the net). See Figure˜1 for an example.
@@
It is easy to observe that $\operatorname{\mathbf{pe}}(S_{k})=\operatorname{\mathbf{pe}}(T_{k})=k$ for each positive integer $k$. Consequently, if $G$ is a graph such that each connected induced subgraph of $G$ has path eccentricity less than $k$, then $G$ is $\{S_{k},T_{k}\}$-free. As the main result of this paper, we show that this necessary condition is in fact also sufficient. We thus obtain the following unifying theorem, valid for all positive integers $k$ (and thus capturing also Corollary˜1.2, which corresponds to the case $k=1$).
@@
The case $k=2$ of Theorem˜1.4 characterizes graphs every connected induced subgraph of which has a dominating path, and implies the following.
@@
Corollary˜1.5 unifies and generalizes several results from the literature regarding sufficient conditions for the existence of a dominating path:
@@
    A graph $G$ is said to be AT-free if it does not have an independent set $I$ such that $|I|=3$ and for every vertex $v\in I$, the remaining two vertices in $I$ are in the same component of GN[v]G-N[v]. Every connected AT-free graph has a dominating path (see Corneil, Olariu, and Stewart [COS97], as well as [COS95]).
@@
    A graph $G$ is biconvex if it is bipartite, with parts $A$ and $B$ that can each be linearly ordered so that for each vertex $v$ of $G$, the neighborhood of $v$ in the part not containing $v$ forms a consecutive segment of vertices with respect to the linear ordering. Gómez and Gutiérrez [GG23] and, independently Antony, Das, Gosavi, Jacob, and Kulamarva [ADG+24], proved that every connected biconvex graph has a dominating path (or, equivalently, a spanning caterpillar).
Indeed, each of the previously stated cases corresponds to a subclass of the class of $\{S_{2},T_{2}\}$-free graphs, hence, Corollary˜1.5 implies all these results.
Further consequences of Theorem˜1.4 are related to a recent work of Bastide, Hilaire, and Robinson [BHR25]. The authors explored a relation between the path eccentricity of a graph and a certain property of its adjacency matrix (or equivalently, on the neighborhoods of its vertices). They introduced the partially augmented consecutive ones property (denoted *-C1P for short): a graph has the *-C1P if there exists an ordering of its vertices sending the open or closed neighborhood of every vertex to a consecutive set. The authors showed that graphs with the *-C1P have path eccentricity at most 2, and conjectured that they actually have path eccentricity at most 1. We show that Theorem˜1.4 resolves the conjecture.
@@
Another result from [BHR25] is related to the following generalization of AT-freeness. A graph $G$ is said to be $k$-AT-free if $G$ does not have an independent set $I$ such that $|I|=3$ and for every vertex $v\in I$, the remaining two vertices in $I$ are in the same component of the graph $G-\{u\colon u$ is at distance at most $k$ of v}v\}. In particular, a graph is AT-free if and only if it is $1$-AT-free. In [BHR25], the authors proved that for every $k\geqslant 1$, if a graph $G$ is $k$-AT-free, then $\operatorname{\mathbf{pe}}(G)\leqslant k$. This result follows from Theorem˜1.4. Indeed, it is easy to check that for every $k\geqslant 2$, the graphs $S_{k}$ and $T_{k-1}$ are not (k1)(k-1)-AT-free. Thus, if a graph $G$ is (k1)(k-1)-AT-free, then $G$ is $\{S_{k},T_{k}\}$-free, which implies by Theorem˜1.4 that $\operatorname{\mathbf{pe}}(G)<k$.
For the case where a single induced subgraph is excluded, Theorem˜1.4 leads to the following characterization. The symbol ++ denotes the disjoint union of graphs, and for an integer $k\geqslant 0$ and a graph $H$, we denote by $kH$ the disjoint union of $k$ copies of $H$.
@@
In the case of $3P_{2}$-free graphs, we obtain a much stronger property. We can find a path with eccentricity at most $1$ that is a longest path in the graph. While the proof of this result implies that we can transform a longest path into a longest path that is dominating in polynomial time, it is NP-hard to find a longest path in $3P_{2}$-free graphs. In fact, determining whether a given graph has a Hamiltonian path is NP-complete in the class of split graphs [Mül96], a subclass of $2P_{2}$-free graphs.
@@
Every connected $3P_{2}$-free graph $G$ has a longest path that is dominating.
@@
This is a very fine line, as this property is completely violated in the case of $\{S_{2},T_{2}\}$-free graphs or even $\{P_{5},4P_{2}\}$-free graphs; see Figure˜2. It also cannot be generalized to higher eccentricity; see for example, a $\{3P_{3},P_{7}+P_{2}\}$-free graph in Figure˜3. Moreover, in a $3P_{2}$-free graph, there might exist a longest path that is not dominating, as depicted in Figure˜4.
Figure 3: An example of a $\{3P_{3},P_{7}+P_{2}\}$-free graph no longest path of which is 22-dominating. The vertices depicted in orange need to belong to every 22-dominating path; however, no longest path contains the top orange vertex.
@@
Note that for $3P_{2}$-free graphs, the algorithm does not necessarily find a longest path that is dominating, unless we are given another longest path on the input. The ideas in the proof of Theorem˜1.8 can be generalized to the setting of Corollary˜1.7, providing us with a polynomial-time algorithm that finds a (k1)(k-1)-dominating path. The algorithm is robust in the sense of Spinrad and Raghavan [RS03]: it works on any input graph $G$, not necessarily $3P_{k}$-free or ($P_{2k+1}+P_{k-1}$)-free, and either correctly finds both of these induced subgraphs, or a path with eccentricity less than $k$. Moreover, the integer $k$ can be part of the input, and the running time is independent of $k$.
To state our next result, we require a generalization of the graphs $S_{k}$ and $T_{k}$ to a broader class of graphs, represented in Figure˜5 (three leftmost figures) and defined as follows. Following the notation introduced in [DDH+24], we denote by $\mathcal{S}$ the class of subdivisions of the graph $K_{1,3}$. The class $\mathcal{T}$ is the class of graphs that can be obtained from three paths of length at least one by selecting one endpoint of each path and adding three edges between those endpoints to create a triangle. The class $\mathcal{M}$ is the class of graphs $H$ that consist of a path $P$ and a vertex $a$, called the center of $H$, such that $a$ is nonadjacent to the endpoints of $P$ and $a$ has at least two neighbors in $P$. Given a graph $H\in\mathcal{S}\cup\mathcal{T}\cup\mathcal{M}$, the extremities of $H$ are the vertices of degree one as well as the center of $H$ in case $H\in\mathcal{M}$. Observe that any $H\in\mathcal{S}\cup\mathcal{T}\cup\mathcal{M}$ has exactly three extremities. Note in particular that for every $k\geqslant 1$, $S_{k}\in\mathcal{S}$, and $T_{k}\in\mathcal{T}$. For every $k\geqslant 1$, we denote by $\mathcal{M}_{k}$ the family of graphs obtained by replacing each extremity of a graph $M\in\mathcal{M}$ by the extremity of a $P_{k}$.
$G\in\mathcal{S}$$\in\mathcal{T}GG\in\mathcal{M}$$G\in\mathcal{M}_{k}$
Figure 5: The three leftmost figures are representations of graphs in $\mathcal{S}$, $\mathcal{T}$, and $\mathcal{M}$, respectively from left to right, where the extremities are represented with white vertices; the rightmost figure represents a graph in $\mathcal{M}_{k}$, for some $k\geqslant 1$, where each dashed blue edge represents a $P_{k}$. In all these figures, solid edges represent edges of the graph, wavy edges represent paths of length at least 11, and dotted edges may be present or not.
We are now ready to state our next theorem.
Theorem 1.9.
There is an algorithm running in time $\mathcal{O}(n^{2}(n+m))$ that, given a connected graph $G$ with $n$ vertices and $m$ edges, and an integer $k\geqslant 1$, finds one of the following:
  • a path with eccentricity less than $k$ in $G$, or
  • an induced subgraph $H$ of $G$ isomorphic to either $S_{k}$, $T_{k}$, or a graph in $\mathcal{M}_{k}$.
Let us notice that it is easy to see that in every case of the second outcome, $H$ contains as an induced subgraph both $3P_{k}$ and $P_{2k+1}+P_{k-1}$, giving the aforementioned algorithmic variant of Corollary˜1.7. Moreover, it also holds that $H$ contains either $S_{k}$ minus a leaf or $T_{k}$ minus a leaf as an induced subgraph, bringing us closer to an algorithmic version of Theorem˜1.4.
@@
In Section˜2, we give an overview of known results on structural domination, including a key result of Bacsó and Tuza [BT12] that we use in the proof of our main theorem (Theorem˜1.4). We prove Theorems˜1.4 and 1.7 in Section˜3. We prove Theorem˜1.6 in Section˜4. We prove Theorems˜1.8 and 1.9 in Section˜5. We conclude the paper with some final remarks and open questions in Section˜6.
@@ 2 Background on Structural Domination
Theorem˜1.4 is closely related to the framework of structural domination. Given a class $\mathcal{D}$ of connected graphs, let us say that a graph $G$ is hereditarily dominated by $\mathcal{D}$ if each connected induced subgraph of $G$ contains a dominating set that induces a graph from $\mathcal{D}$. Bacsó, Michalak, and Tuza (see [BMT05]) gave forbidden induced subgraph characterizations of graphs hereditarily dominated by classes of complete bipartite graphs, stars, connected bipartite graphs, and complete $k$-partite graphs. Similar results were obtained by Michalak (see [Mic07]) for graphs hereditarily dominated by classes of cycles and paths, paths, trees, and trees with bounded diameter. Camby and Schaudt (see [CS16]) proved that for every $k\geqslant 4$, a graph $G$ is $P_{k}$-free if and only if it is hereditarily dominated by the class consisting of all connected $P_{k-2}$-free graphs along with the $k$-vertex cycle.
The most general result along these lines was obtained independently by Bacsó and Tuza (see [Tuz08, Bac09]). To state the result, we need one more definition. Given a connected graph $H$, the leaf graph of $H$ is the graph $H^{\text{\textleaf}}$ obtained from $H$ by attaching a leaf to each vertex $v\in V(H)$ that is not a cut-vertex (that is, the graph HvH-v is connected). By definition, these new leaves are pairwise distinct and non-adjacent.
@@
The main result of [Tuz08, Bac09] is even more general, without the restriction that $\mathcal{D}$ contains the class of all paths; but since we do not need this result, we omit the precise statement. A generalization of Theorem˜2.1 to domination at distance was proved by Bacsó and Tuza (see [BT12]). A distance-$k$ dominating set in a graph $G$ is a set $D\subseteq V(G)$ such that every vertex in $G$ is at distance at most $k$ from some vertex in $D$. If, in addition, the graph $G[D]$ is connected, then $D$ is said to be a connected distance-$k$ dominating set. A graph $G$ is said to be hereditarily $k$-dominated by a class $\mathcal{D}$ of connected graphs if each connected induced subgraph of $G$ admits a distance-$k$ dominating set that induces a graph from $\mathcal{D}$. Given a connected graph $H$, the $k$-leaf graph of $H$ is the graph $H_{k}^{\text{\textleaf}}$ obtained from $H$ by attaching a pendant path of length $k$ to each vertex $v\in V(H)$ that is not a cut-vertex. Bacsó and Tuza (see [BT12]) generalized Theorem˜2.1 as follows.
@@
Note that Theorem˜2.1 is a special case of Theorem˜2.2 for the case $k=1$.
3 Proof of General Theorem (Theorem˜1.4)
@@
Fix an integer $k\geqslant 1$. Let $G$ be a graph such that every connected induced subgraph $H$ of $G$ satisfies $\operatorname{\mathbf{pe}}(H)<k$. To show that $G$ is $\{S_{k},T_{k}\}$-free, it suffices to show that $\operatorname{\mathbf{pe}}(S_{k})\geqslant k$ and $\operatorname{\mathbf{pe}}(T_{k})\geqslant k$. Let $H$ be a graph isomorphic to either $S_{k}$ or $T_{k}$ and let $S$ be any path in $H$. We claim that $S$ has eccentricity at least $k$. Suppose not. Let $C$ be the set of vertices of degree $3$ in $H$. Then, the graph HV(C)H-V(C) consists of three disjoint paths $P$, $Q$, and $R$, each with $k$ vertices. Moreover, each of the paths $P$, $Q$, and $R$ contains a vertex at distance $k$ from $V(C)$. Since the eccentricity of $S$ is less than $k$, each of the paths $P$, $Q$, and $R$ contains a vertex of $S$. Since $S$ has at most two endpoints, at most two of the paths $P$, $Q$, and $R$ contain an endpoint of $P$; hence, we may assume without loss of generality that $P$ contains an internal vertex of $S$ but no endpoint of $S$. However, this implies that two distinct edges of $S$ connect $P$ with the rest of $H$, a contradiction. This shows that $\operatorname{\mathbf{pe}}(H)\geqslant k$, as claimed.
It remains to prove that if $G$ is $\{S_{k},T_{k}\}$-free, then every connected induced subgraph $H$ of $G$ has path eccentricity less than $k$, or, equivalently, that every connected $\{S_{k},T_{k}\}$-free graph has path eccentricity less than $k$. For $k=1$, the statement coincides with that of Theorem˜1.1. Suppose now that $k\geqslant 2$. Let $\mathcal{F}=\{S_{1},T_{1}\}$ and let $\mathcal{D}$ be the class of all connected $\mathcal{F}$-free graphs. Note that $\mathcal{F}_{k-1}^{\text{\textleaf}}=\{S_{k},T_{k}\}$. Consequently, since $\mathcal{D}$ satisfies the hypothesis of Theorem˜2.2, it follows that every $\{S_{k},T_{k}\}$-free graph $G$ is hereditarily (k1)(k-1)-dominated by $\mathcal{D}$. Let $H$ be a connected induced subgraph of $G$. Then, since $G$ is hereditarily (k1)(k-1)-dominated by $\mathcal{D}$, there exists a distance-(k1)(k-1) dominating set $D$ in $H$ such that the induced subgraph $H[D]$ belongs to $\mathcal{D}$. By Theorem˜1.1, the graph $H[D]$ admits a Hamiltonian path $P$. It follows that $P$ is a path in $H$ with eccentricity less than $k$, implying that $\operatorname{\mathbf{pe}}(H)<k$. ∎
Our proof of Corollary˜1.7 will be based on the following.